OpenStudy (anonymous):

find the reciprocal 7/2a

OpenStudy (221emily):

Do you know how to find reciprocals and you are stuck on this one? or do you have no clue what and how to come up with reciprocals?

OpenStudy (anonymous):

i just have no clue on what the reciprocals for 7/2a is.

OpenStudy (221emily):

ok just a second let me draw something

OpenStudy (anonymous):

alright

OpenStudy (221emily):

the reciprocal of \[\frac{ 6 }{ 5 }\] is \[\frac{ 5 }{ 6 }\] right?

OpenStudy (anonymous):

right.

OpenStudy (221emily):

so does your problem look like this\[\frac{ 7 }{ 2a }\] ?

OpenStudy (anonymous):

yes

OpenStudy (221emily):

What did I do to the first fraction to get the second one?

OpenStudy (anonymous):

you flipped the denominator & the numerator.

OpenStudy (221emily):

Yes!! (continued)

OpenStudy (221emily):

\[\frac{ 5z }{ 6d }\] and \[\frac{ 6d }{ 5z }\] right?

OpenStudy (anonymous):

right ..i think i get it now thanks :)

OpenStudy (221emily):

yep anytime :)